OpenStudy (anonymous):

Two families paid a total of $55.00 to see a movie. Each adult ticket cost $7.50. If there were 3 adults and 5 children, how much did each child's ticket cost? $5.50 $6.00 $6.50 $7.00

OpenStudy (cwrw238):

3*7.5 + 5*x = 55 where x is cost of childs ticket solve this for x

OpenStudy (cwrw238):

can you do this?

OpenStudy (anonymous):

not really

OpenStudy (cwrw238):

ok - first simplify 22.5 + 5x = 55 subtract 22.5 from both sides to get 5x on its own 5x = 55-22.5 5x = 32.5 now divide both sides by 5 x = 32.5 / 5 = 6.5 so its $6.50

OpenStudy (anonymous):

thanks